OpenStudy (anonymous):

u=-4i+1j and v=4i+1j find ||u+v||.

OpenStudy (tkhunny):

Have you considered adding the two vectors?

OpenStudy (anonymous):

just out of curiosity the i in the problem doesnt connect to the i symbol used in algebra does it

OpenStudy (anonymous):

and what do you mean

OpenStudy (tkhunny):

'i' indicates one direction. 'j' indicates an orthogonal direction. Just add them. How else can you calculate the magnitude of their sum?

OpenStudy (anonymous):

im not understnding can you explain how i add them? there just letters as far as i can tell they have no given value

OpenStudy (anonymous):

wait its not that easy is it???

OpenStudy (anonymous):

i just add -4i and 1j?

OpenStudy (tkhunny):

Add these: (4dogs+ 2cats) + (1dog+3cats)

OpenStudy (anonymous):

5 dogs and 5 cats

OpenStudy (tkhunny):

Perfect. Do that with 'i' and 'j'.

OpenStudy (anonymous):

so my answer is u+v=1i and 2j

OpenStudy (tkhunny):

How did you get 1i? -4i + 4i = ??

OpenStudy (anonymous):

-4+4

OpenStudy (anonymous):

or is it 0 so u+v=2j

OpenStudy (tkhunny):

That's right.Since the 'i's went away, the magnitude is now trivial.

OpenStudy (anonymous):

ohhh my bad it just seems to easy for a pre calc problem

OpenStudy (tkhunny):

Had there been a non-zero 'i', you would have needed the Distance Formula.

OpenStudy (anonymous):

non zero i?

OpenStudy (anonymous):

and also what are the lines on either side of u+v for

OpenStudy (tkhunny):

(i + 2j)+(i +j) = 2i + 3j NonZero iand nonzero j

OpenStudy (anonymous):

ohh

OpenStudy (tkhunny):

The lines indicate magnitude. It's value can be calculated with the distance formula.

OpenStudy (anonymous):

do i need to do that

OpenStudy (anonymous):

ohhh wait there was one other part of the problem i didnt see

OpenStudy (anonymous):

Find the specified vector or scalar

OpenStudy (anonymous):

does that mean anything else needs to be done?

OpenStudy (tkhunny):

The sum is 2j- a vector. The magnitude is 2 - a scalar.

OpenStudy (anonymous):

what does that mean and how do you get that

OpenStudy (tkhunny):

We just went through it all. Go reread until it soaks in.

OpenStudy (anonymous):

what is a vector and a scalr is what i mean